# Advancing GPU IPC for Stiff Affine-Deformable Simulation

### v2

@@ 7.1. Ablation and Comparative Studies
As seen from the similar average CG iterations across resolutions (the rows for Figure 11 in Table 1), our preconditioner maintains the scalability of MAS, resulting in near-linear total timing with respect to resolution. In Figure 13, the ’Ours (16)’ configuration shows only a 1.17$\times$ speedup in convergence, as the volumetric structure of the bunnies results in similar METIS and Morton code reorderings, limiting improvement. However, we still observe a 1.58$\times$ speedup in PCG time and a 1.47$\times$ overall speedup, due to the lower computational cost of our preconditioner compared to MAS.
Table 1. Simulation statistics with different preconditioners. Mesh DOFs: number of vertices (v), tetrahedra (t), and surface triangles (f). PPC, PPR, and CGR represent the time spent on preconditioner pre-computation, preconditioning, and CG iterations, respectively, with PCG = PPC + PPR + CGR. The remaining columns report the time for miscellaneous tasks (misc), total simulation time (timeTot), total CG iterations (#cg), total Newton iterations (#Newton), and the average CG iterations per Newton iteration (avg. #cg). All times are in seconds.
Example Mesh DOFs: v, t, f Preconditioner PPC PPR CGR PCG Misc TimeTot #cg #Newton avg. #cg
Figure 11 $42K,171K,45K$ GPU MAS 38 55 102 195 26 221 6.98e5 2.03e3 344
Block Diagonal 0.06 21 338 359 27 385 2.24e6 2.07e3 1082
Ours (32) 32 25 60 117 26 143 3.45e5 2.02e3 171
Ours (16) 10 29 69 108 26 134 4.41e5 1.99e3 221
@@
Figure 11 $114K,558K,69K$ GPU MAS 105 162 257 524 69 593 9.51e5 2.31e3 412
Block Diagonal 0.16 43 582 625 69 694 2.48e6 2.30e3 1079
Ours (32) 96 79 127 302 69 371 4.37e5 2.31e3 189
Ours (16) 33 60 145 238 69 307 5.20e5 2.31e3 225
@@
Figure 13 $203K,1122K,50K$ GPU MAS 152 188 229 569 87 656 6.23e5 1.76e3 354
Block Diagonal 0.18 63 918 981 88 1069 2.61e6 1.77e3 1474
Ours (32) 134 137 180 451 87 538 4.54e5 1.76e3 258
Ours (16) 42 108 210 360 87 447 5.30e5 1.76e3 301
@@
Figure 19 $123K,330K,211K$ GPU MAS 371 836 1029 2236 217 2453 4.38e6 7.93e3 552
Block Diagonal 0.44 137 1920 2057 216 2273 8.84e6 7.90e3 1119
Ours (32) 354 196 298 848 215 1063 9.98e5 7.86e3 127
Ours (16) 76 215 439 730 215 945 1.81e6 7.85e3 231
@@
Figure 20 $79K,298K,110K$ GPU MAS 86 193 274 553 42 595 1.42e6 1.62e3 877
Block Diagonal 0.08 57 798 855 43 898 5.09e6 1.64e3 3105
Ours (32) 47 89 124 260 42 302 6.84e5 1.63e3 420
Ours (16) 12 68 136 216 41 257 7.53e5 1.62e3 465
@@
Figure 12 presents a cloth simulation example using FBW (Kim, 2020) with and without barrier-based (Li et al., 2021) or our cubic strain-limiting energy. The cloth consists of 100K nodes and is tested with various membrane stiffnesses. In cloth simulation, avoiding excessive stretching is crucial, as it can cause unrealistic, overly-elastic behavior. To prevent this, a large Young’s modulus (e.g., $5MPa$), within the range of real cloth parameters (Penava et al., 2014), is needed. However, such stiffness can lead to membrane locking in even medium-resolution meshes, resulting in sharp creases and artificial plastic appearances (12(a)). Reducing the stiffness to $0.5MPa$ mitigates membrane locking but visible artifacts remain (12(b)). Further reducing the stiffness to $50KPa$ smooths wrinkles but causes over-elongation (12(c)). Our cubic strain-limiting energy provides sufficient stretching resistance as in real fabrics, avoiding excessive stretching and capturing smooth, realistic wrinkles with a small membrane stiffness of $50KPa$ for compression (12(e), and supplemental video).
We also compare with the barrier-based strain-limiting method in CIPC using their default settings: a $1.1$ strain limit and $1$ activation threshold. We use a $1KPa$ stiffness for the barrier term, as smaller or larger values both lead to worse conditioned systems. With similar result quality (12(d)), our method is 1.24×1.24$\times$1.24 × faster (Table 2), with lower costs in local Hessian construction and line search (see Table 3 for a more detailed breakdown). While our method performs well for practical cloth manipulation scenarios, it does not guarantee strain limit satisfaction in extreme cases like the barrier-based method. Thus, we further compare their behavior in a stress test (Figure 22), detailed later.
@@
As illustrated in Figure 14, we use the wrecking ball simulation (Figure 18) to conduct an ablation study on our multi-layer reduction strategy (Strategy C) for global Hessian matrix assembly, comparing it against single-layer reduction (Strategy B) and pure atomic operations (Strategy A). In both Strategies A and B, the local Hessian matrices of all contact pairs are first mapped to the affine body DOFs by multiplying the Jacobian matrices. Strategy B then performs a reduction to accumulate these matrices into the global Hessian, while Strategy A uses atomic operations directly for accumulation. To ensure a fair comparison, we also sort the value array prior to atomic accumulation. This step optimizes the distribution of the value array, enabling aggregation optimization in the CUDA intrinsic function atomicAdd, thereby achieving the best possible performance. Even with this optimization, pure atomic operations remain slightly slower than single-layer reduction, as shown in the comparison of Strategies A and B in Figure 14. Our multi-layer reduction strategy achieves an additional nearly 5×5$\times$5 × speedup over single-layer reduction.
@@
We compare our FastHashReduction method with the CUB intrinsic function cub::DeviceReduce::ReduceByKey and plot our speedup relative to the number of values being accumulated (Figure 15). To keep the total usage of GPU memory unchanged, the reduction is performed on 16M $3\times 3$ matrices, and we adjust the number of matrices with the same row or column indices to control the workload. Our method shows superior performance across a wide range of accumulation workload, from $1$ to $2^{16}$, achieving a maximum speedup of approximately 1.55×1.$5$ at $2^{11}$. In FEM simulations, the average number of values with the same row or column indices typically remains below $2^{7}$, where our method achieves a consistent speedup of $1.2$ to 1.4×1.4$\times$1.4 ×. Additionally, we observe that cub::DeviceReduce::ReduceByKey fails to compile when reducing the $12\times 12$ affine body Hessian matrix due to shared memory limitations. In contrast, our FastHashReduction method remains effective by reusing warp registers, leveraging the element-independent nature of matrix summation.
@@
We compare the performance of different SpMV methods using the stiff and soft bunnies example shown in Figure 13, simulating both bunnies with FEM. As shown in Figure 16, the BSR SpMV (1$\times$) from cuSparse [^1] https://developer.nvidia.com/cusparse is treated as a baseline. Without matrix sorting, the Triplet SpMV (0.16$\times$) implemented by atomically adding the matrix-vector products and the MatrixFree SpMV (0.18$\times$) from GIPC both suffer from an atomic operation flood, which undermines the parallelism. On the other hand, sorting-based methods have a significant speedup, benefiting from the reduction of atomic operations. The traditional CSR SpMV (0.75$\times$) from cuSparse has an almost 5$\times$ speedup compared to the unsorted methods. The Block Coordinates (BCOO) SpMV from MUDA [^2] https://github.com/MuGdxy/muda and the BSR SpMV from cuSparse further improves the performance by applying blockwise matrix-vector multiplication. Our SRBK SpMV method (1.85×1.85$\times$1.85 ×) further mitigates the memory-bound issue of BSR SpMV and boosts the performance by almost 2×2$\times$2 ×.
To evaluate the scalability of our method, we compare its speedup factor to BSR on matrices with varying numbers of non-zero $3\times 3$ blocks per row. As shown in Figure 17, the speedup factor increases from 0.93×0.93$\times$0.93 × to 1.9×1.9$\times$1.9 × as the number of non-zero blocks approaches $2^{3}$ from 0. For average non-zero block counts between $2^{3}$ and $2^{6}$ – the typical range for FEM global systems – the speedup factor fluctuates between 1.8×1.8$\times$1.8 × and 2×2$\times$2 ×, which demonstrates strong scalability.
Figure 14. Ablation study on multi-layer reduction for Hessian assembly. We compare our multi-layer reduction strategy (C) against single-layer reduction (B) and pure atomic operations (A) in the wrecking ball simulation (Figure 18). Despite sorting the value array for optimal atomicAdd performance, using atomic operations is slightly slower than single-layer reduction, and ours achieves nearly 5×5$\times$5 × speedup over Strategy B.
Figure 15. Speedup of FastHashReduction vs. CUB ReduceByKey. Speedup factor is shown relative to the average number of matrices sharing the same row or column indices. A peak speedup of 1.55×1.$5$ occurs around $2^{11}$. For typical duplication levels in FEM simulations (up to $2^{7}$), FastHashReduction maintains a consistent speedup between 1.2×1.2$\times$1.2 × and 1.4×1.4$\times$1.4 ×.
@@
Table 3. Simulation statistics for all examples. Columns show time for energy gradient and Hessian computation (Hess), linear system solves (LSolve), CCD (CCD), backtracking line search (LineS), miscellaneous tasks (Misc), and the entire simulation (TimeTot), followed by iteration counts, and finally, speedup relative to GIPC. All times are in seconds. Here, we compare our method with GIPC (Huang et al., 2024), ZeMa (Du et al., 2024), and the barrier-based SL (Li et al., 2021). We also show how each of our innovation contributes to the overall speedup.
@@
Mesh DOFs: v, t, f Framework Hess LSolver CCD LineS Misc TimeTot #cg #Newton avg. #cg Speed Up
Figure 11 114K, 558K, 69K GIPC 63 1134 5 13 6 1221 9.41e5 2.30e3 409 1.00$\times$
114K, 558K, 69K Ours (SRBK) 45 524 5 13 6 593 9.51e5 2.31e3 412 2.06$\times$
114K, 558K, 69K Ours (CEMAS+SRBK) 45 238 5 13 6 307 5.20e5 2.31e3 225 3.98$\times$
Figure 12 109K, 80K, 200K GIPC 106 606 20 54 6 792 2.34e5 3.89e3 60 1.00$\times$
109K, 80K, 200K Ours (SRBK) 75 489 20 55 6 645 2.37e5 3.86e3 61 1.23$\times$
109K, 80K, 200K Ours (CEMAS+SRBK) 75 184 20 55 6 340 2.61e5 3.95e3 66 2.33$\times$
109K, 80K, 200K CIPC (CEMAS+SRBK) 104 202 21 91 6 424 2.89e5 4.20e3 69 -
100K, 0, 200K Ours (ABD+CEMAS+SRBK) 67 145 19 51 6 288 2.06e5 3.76e3 55 2.75$\times$
Figure 13 203K, 1122K, 50K GIPC 90 1481 4 14 7 1596 6.23e5 1.76e3 354 1.00$\times$
114K, 561K, 50K ZeMa 42 54810 2.7 9 7 54871 - 1.74e3 - 0.03$\times$
203K, 1122K, 50K Ours (SRBK) 63 569 3 14 7 656 6.23e5 1.76e3 354 2.43$\times$
203K, 1122K, 50K Ours (CEMAS+SRBK) 63 360 3 14 7 447 5.30e5 1.76e3 301 3.57$\times$
114K, 561K, 50K Ours (ABD+CEMAS+SRBK) 37 104 2.7 8.5 7 159 1.73e5 1.73e3 100 10.03$\times$
Figure 19 123K, 330K, 211K GIPC 137 3910 35 83 8 4173 4.31e6 7.87e3 548 1.00$\times$
123K, 330K, 211K Ours (SRBK) 91 2236 35 83 8 2453 4.38e6 7.93e3 552 1.70$\times$
123K, 330K, 211K Ours (CEMAS+SRBK) 90 730 35 82 8 945 1.81e6 7.85e3 231 4.42$\times$
Figure 20 79K, 298K, 110K GIPC 24 1080 5 13 8 1130 1.42e6 1.62e3 877 1.00×1.$0$
79K, 298K, 110K Ours (SRBK) 16 553 5 13 8 595 1.42e6 1.62e3 877 1.90$\times$
79K, 298K, 110K Ours (CEMAS+SRBK) 16 216 5 12 8 257 7.53e5 1.62e3 465 4.40$\times$
78K, 290K, 110K Ours (ABD+CEMAS+SRBK) 14 148 4 8 7 181 4.50e5 1.41e3 319 6.24$\times$
Figure 21 70K, 241K, 106K GIPC 692 10686 114 637 14 12143 7.30e6 1.58e4 462 1.00$\times$
70K, 241K, 106K Ours (SRBK) 460 4511 113 594 14 5692 7.58e6 1.58e4 480 2.13$\times$
70K, 241K, 106K Ours (CEMAS+SRBK) 452 1547 112 597 14 2722 3.37e6 1.57e4 214 4.46$\times$
53K, 133K, 106K Ours (ABD+CEMAS+SRBK) 407 696 103 535 12 1753 1.48e6 1.46e4 101 6.93$\times$
Figure 22 105K, 10K, 202K GIPC 210 2473 54 211 9 2957 3.10e6 9.43e3 329 1.00$\times$
105K, 10K, 202K Ours (SRBK) 142 762 51 205 9 1069 3.13e6 9.27e3 338 2.76$\times$
105K, 10K, 202K Ours (CEMAS+SRBK) 145 451 52 207 9 864 8.21e5 9.50e3 86 3.42$\times$
105K, 10K, 202K CIPC (CEMAS+SRBK) 219 812 59 329 10 1429 1.75e6 9.69e3 180 -
105K, 0, 202K Ours (ABD+CEMAS+SRBK) 161 468 47 178 9 863 7.78e5 8.51e3 91 3.42$\times$
@@
Here, we compare our method with GIPC (Huang et al., 2024) and ZeMa (Du et al., 2024) on the stiff and soft bunnies example (Figure 13). In this simulation, the Young’s modulus is set to $10^{6}$ Pa for the bottom bunny and $10^{8}$ Pa for the top bunny. Our framework can treat the top bunny as a stiff affine body and the bottom bunny with FEM, while GIPC needs to simulate both using FEM. Results in Table 3 indicate that our coupling framework achieves a 10$\times$ overall speedup compared to GIPC. To analyze the sources of this acceleration, we applied our contributions incrementally. ’Ours (SRBK)’ uses only our SRBK method, matching other components to GIPC, which already yields a 2.43×2.43$\times$2.43 × speedup with more efficient Hessian assembly and SpMV. ’Ours (CEMAS+SRBK)’ further adds connectivity-enhanced (CE) MAS preconditioner to ’Ours (SRBK)’, retaining other components, and shows an additional 1.47×1.47$\times$1.47 × overall speedup due to faster PCG convergence enabled by CEMAS, resulting in a 3.57×3.57$\times$3.57 × improvement compared to GIPC. Note that in both ’Ours (SRBK)’ and ’Ours (CEMAS+SRBK)’ cases, we simulate both bunnies using FEM. ZeMa performs the worst in this case, as direct solvers are not efficient for simulations with a large number of DOFs. Thus, for general IPC simulations, GIPC remains the best alternative. Therefore, in subsequent experiments, we compare our methods only against GIPC.
@@
We further compare our method with GIPC on the examples shown in Figure 11, Figure 12, Figure 19, Figure 20, Figure 22, and Figure 21, with statistics reported in Table 3. Generally, ’Ours (SRBK)’ achieves a 1.23×1.23$\times$1.23 ×2.76×2.76$\times$2.76 × speedup over GIPC. Adding our CEMAS preconditioner further improves performance, providing an additional 1.24×1.24$\times$1.24 ×2.6×2.6$\times$2.6 × speedup as shown by ’Ours (CEMAS+SRBK)’, leading to an overall improvement of 2.33×2.$3$4.46×4.46$\times$4.46 × compared to GIPC. Notably, this is achieved without approximating stiff materials as affine bodies, which can further boost performance up to 2.75×2.75$\times$2.75 ×10×10$\times$10 × faster than GIPC. In Figure 12, comparing ’Ours (SRBK)’ and ’Ours (CEMAS+SRBK)’ reveals that while CEMAS does not improve PCG convergence due to similar ordering with Morton code sorting in GIPC, it significantly reduces PCG solver time due to lower construction and preconditioning costs. Here, PCG costs represent a smaller portion of total timing than in other examples due to fewer PCG iterations required, resulting in a modest 1.23×1.23$\times$1.23 × speedup with SRBK alone. We will discuss Figure 21 and Figure 22 in detail in the next subsection on stress tests.

### v3

@@ 7.1. Ablation and Comparative Studies
As seen from the similar average CG iterations across resolutions (the rows for Figure 11 in Table 1), our preconditioner maintains the scalability of MAS, resulting in near-linear total timing with respect to resolution. In Figure 14, the ’Ours (16)’ configuration shows only a 1.17$\times$ speedup in convergence, as the volumetric structure of the bunnies results in similar METIS and Morton code reorderings, limiting improvement. However, we still observe a 1.58$\times$ speedup in PCG time and a 1.47$\times$ overall speedup, due to the lower computational cost of our preconditioner compared to MAS. Notably, as the number of bunnies increases or their spatial distribution becomes more irregular, our preconditioner achieves greater speedup (up to 3.5$\times$ in PCG time). We also tested a heterogeneous scene with 9 bunnies, where the Young’s modulus varies from 20 KPa to 10 MPa, and observed consistent PCG solver speedup compared to the homogeneous counterpart. Additionally, in a sequence of simulations with an increasing number of contact pairs between two plates, our preconditioner also maintained consistent speedup. These additional experiments and analyses are provided in the supplemental document.
Table 1. Simulation statistics with different preconditioners. Mesh DOFs: number of vertices (v), tetrahedra (t), and surface triangles (f). PPC, PPR, and CGR represent the time spent on preconditioner pre-computation, preconditioning, and CG iterations, respectively, with PCG = PPC + PPR + CGR. The remaining columns report the time for miscellaneous tasks (misc), total simulation time (timeTot), total CG iterations (#cg), the average Newton iterations per frame (avg. #Newton), and the average CG iterations per Newton iteration (avg. #cg). All times are in seconds.
Example Mesh DOFs: v, t, f Preconditioner PPC PPR CGR PCG Misc TimeTot #cg avg. #Newton avg. #cg
Figure 11 $42K,171K,45K$ GPU MAS 38 55 102 195 26 221 6.98e5 11.28 344
Block Diagonal 0.06 21 338 359 27 385 2.24e6 11.50 1082
Ours (32) 32 25 60 117 26 143 3.45e5 11.22 171
Ours (16) 10 29 69 108 26 134 4.41e5 11.06 221
@@
Figure 11 $114K,558K,69K$ GPU MAS 105 162 257 524 69 593 9.51e5 12.78 412
Block Diagonal 0.16 43 582 625 69 694 2.48e6 12.78 1079
Ours (32) 96 79 127 302 69 371 4.37e5 12.83 189
Ours (16) 33 60 145 238 69 307 5.20e5 12.83 225
@@
Figure 14 $203K,1122K,50K$ GPU MAS 152 188 229 569 87 656 6.23e5 9.78 354
Block Diagonal 0.18 63 918 981 88 1069 2.61e6 9.83 1474
Ours (32) 134 137 180 451 87 538 4.54e5 9.78 258
Ours (16) 42 108 210 360 87 447 5.30e5 9.78 301
@@
Figure 20 $123K,330K,211K$ GPU MAS 371 836 1029 2236 217 2453 4.38e6 39.65 552
Block Diagonal 0.44 137 1920 2057 216 2273 8.84e6 39.50 1119
Ours (32) 354 196 298 848 215 1063 9.98e5 39.30 127
Ours (16) 76 215 439 730 215 945 1.81e6 39.25 231
@@
Figure 21 $79K,298K,110K$ GPU MAS 86 193 274 553 42 595 1.42e6 8.10 877
Block Diagonal 0.08 57 798 855 43 898 5.09e6 8.20 3105
Ours (32) 47 89 124 260 42 302 6.84e5 8.15 420
Ours (16) 12 68 136 216 41 257 7.53e5 8.10 465
@@
Figure 12 presents a cloth simulation example using FBW (Kim, 2020) with and without barrier-based (Li et al., 2021) or our cubic inexact strain-limiting energy. The cloth consists of 100K nodes and is tested with various membrane stiffnesses. In cloth simulation, avoiding excessive stretching is crucial, as it can cause unrealistic, overly-elastic behavior. To prevent this, a large Young’s modulus (e.g., $5MPa$), within the range of real cloth parameters (Penava et al., 2014), is needed. However, such stiffness can lead to membrane locking in even medium-resolution meshes, resulting in sharp creases and artificial plastic appearances (12(a)). Reducing the stiffness to $0.5MPa$ mitigates membrane locking but visible artifacts remain (12(b)). Further reducing the stiffness to $50KPa$ smooths wrinkles but causes over-elongation (12(c)). Our cubic inexact strain-limiting energy provides sufficient stretching resistance as in real fabrics, avoiding excessive stretching and capturing smooth, realistic wrinkles with a small membrane stiffness of $50KPa$ for compression (12(e), and supplemental video).
We also compare with the barrier-based strain-limiting method in CIPC using their default settings: a $1.1$ strain limit and $1$ activation threshold. We use a $1KPa$ stiffness for the barrier term, as smaller or larger values both lead to worse conditioned systems. With similar result quality (12(d)), our method is 1.24×1.24$\times$1.24 × faster (Table 2), with lower costs in local Hessian construction and line search (see Table 3 for a more detailed breakdown). While our method performs well for practical cloth manipulation scenarios, it does not guarantee strain limit satisfaction in extreme cases like the barrier-based method.
To quantitatively assess constraint satisfaction, we conducted an experiment by pinning the upper corners of a cloth and letting it hang under gravity (Figure 13). The upper corner triangle experienced the most significant stretching. In this setup, the log barrier strain-limiting method strictly constrained the principal stretch to the predefined upper bound of 1.1, with a maximum of 1.098 and an average of 1.012. In contrast, our method with default stiffness (5 MPa) resulted in a maximum principal stretch of 1.35 while maintaining an average of 1.010. Although it does not strictly enforce strain limits at the upper corners, the overall simulation remains physically plausible (see the supplementary video). For stress testing purposes, increasing our stiffness to 0.5 GPa brings the maximum principal stretch in line with the log barrier method, achieving a maximum of 1.100 and an average of 1.001. Furthermore, we also compare our inexact strain limiting with the log barrier method in another stress test (see Figure 23), which is discussed in detail later.
Figure 13. Quantitative evaluation of principal stretch. We simulate a hanging cloth with pinned upper corners. The blue cloth uses the log barrier strain-limiting method (Li et al., 2021) with their default stiffness of 1 KPa and strain limit of 1.1. The red cloth applies our inexact strain-limiting method with a default stiffness of 5 MPa, resulting in a maximum principal stretch of 1.35 at the corners and an average of 1.01. The yellow cloth also uses our method but with an increased stiffness of 500 MPa, reducing the maximum stretch to 1.1 and the average to 1.001.
@@
As illustrated in Figure 15, we use the wrecking ball simulation (Figure 19) to conduct an ablation study on our two-level reduction strategy (Strategy C) for global Hessian matrix assembly, comparing it against single-level reduction (Strategy B) and pure atomic operations (Strategy A). In both Strategies A and B, the local Hessian matrices of all contact pairs are first mapped to the affine body DOFs by multiplying the Jacobian matrices. Strategy B then performs a reduction to accumulate these matrices into the global Hessian, while Strategy A uses atomic operations directly for accumulation. To ensure a fair comparison, we also sort the value array prior to atomic accumulation. This step optimizes the distribution of the value array, enabling aggregation optimization in the CUDA intrinsic function atomicAdd, thereby achieving the best possible performance. Even with this optimization, pure atomic operations remain slightly slower than single-level reduction, as shown in the comparison of Strategies A and B in Figure 15. Our two-level reduction strategy achieves an additional nearly 5×5$\times$5 × speedup over single-level reduction.
@@
We compare our FastHashReduction method with the CUB intrinsic function cub::DeviceReduce::ReduceByKey and plot our speedup relative to the number of values being accumulated (Figure 16). To keep GPU memory usage constant, the reduction is performed on 16M $3\times 3$ matrices. The workload is controlled by adjusting the average number of matrices sharing the same row and column indices, ranging from $1$ to $2^{16}$. Our method shows superior performance across a wide range of accumulation workloads, achieving a maximum speedup of approximately 1.55×1.$5$ at $2^{11}$. In FEM simulations, the average number of values with the same row and column indices typically remains below $2^{7}$, where our method achieves a consistent speedup of $1.2$ to 1.4×1.4$\times$1.4 ×. Additionally, we observe that cub::DeviceReduce::ReduceByKey fails to compile when reducing the $12\times 12$ affine body Hessian matrix due to shared memory limitations. In contrast, our FastHashReduction method remains effective by reusing warp registers, leveraging the element-independent nature of matrix summation.
@@
We compare the performance of different SpMV methods using the stiff and soft bunnies example shown in Figure 14, simulating both bunnies with FEM. As shown in Figure 17, the BSR SpMV (1$\times$) from cuSparse [^3] https://developer.nvidia.com/cusparse is treated as a baseline. Without matrix sorting, the Triplet SpMV (0.16$\times$) implemented by atomically adding the matrix-vector products and the MatrixFree SpMV (0.18$\times$) from GIPC both suffer from an atomic operation flood, which undermines the parallelism. On the other hand, sorting-based methods have a significant speedup, benefiting from the reduction of atomic operations. The traditional CSR SpMV (0.75$\times$) from cuSparse has an almost 5$\times$ speedup compared to the unsorted methods. The Block Coordinates (BCOO) SpMV from MUDA [^4] https://github.com/MuGdxy/muda and the BSR SpMV from cuSparse further improves the performance by applying blockwise matrix-vector multiplication. Our SRBK SpMV method (1.85×1.85$\times$1.85 ×) further mitigates the memory-bound issue of BSR SpMV and boosts the performance by almost 2×2$\times$2 ×. Here, RBK refers to our Reduce-By-Key approach that loads the full Hessian matrix from global memory without leveraging the symmetric structure, leading to degraded performance.
To evaluate the scalability of our method, we compare its speedup factor to BSR on matrices with varying numbers of non-zero $3\times 3$ blocks per row. As shown in Figure 18, the speedup factor increases from 0.93×0.93$\times$0.93 × to 1.9×1.9$\times$1.9 × as the number of non-zero blocks approaches $2^{3}$ from 0. For average non-zero block counts between $2^{3}$ and $2^{6}$ – the typical range for FEM global systems – the speedup factor fluctuates between 1.8×1.8$\times$1.8 × and 2×2$\times$2 ×, which demonstrates improved scalability.
Figure 15. Ablation study on two-level reduction for Hessian assembly. We compare our two-level reduction strategy (C) against single-level reduction (B) and pure atomic operations (A) in the wrecking ball simulation (Figure 19). Despite sorting the value array for optimal atomicAdd performance, using atomic operations is slightly slower than single-level reduction, and ours achieves nearly 5×5$\times$5 × speedup over Strategy B.
Figure 16. Speedup of FastHashReduction vs. CUB ReduceByKey. Speedup factor is shown relative to the average number of matrices sharing the same row and column indices. A peak speedup of 1.55×1.$5$ occurs around $2^{11}$. For typical duplication levels in FEM simulations (up to $2^{7}$), FastHashReduction maintains a consistent speedup between 1.2×1.2$\times$1.2 × and 1.4×1.4$\times$1.4 ×.
@@
Table 3. Simulation statistics for our examples. Columns show time for energy gradient and Hessian computation (Hess), linear system solves (LSolve), CCD (CCD), backtracking line search (LineS), miscellaneous tasks (Misc), and the entire simulation (TimeTot), followed by iteration counts, and finally, speedup relative to GIPC. All times are in seconds. Here, we compare our method with GIPC (Huang et al., 2024), ZeMa (Du et al., 2024), and the barrier-based SL (Li et al., 2021). We also show how each of our innovation contributes to the overall speedup.
@@
# verts, tets, faces Framework Hess LSolver CCD LineS Misc TimeTot #cg avg. #Newton avg. #cg Speed Up
Figure 11 114K, 558K, 69K GIPC 63 1134 5 13 6 1221 9.41e5 12.78 409 1.00$\times$
114K, 558K, 69K Ours (SRBK) 45 524 5 13 6 593 9.51e5 12.83 412 2.06$\times$
114K, 558K, 69K Ours (CEMAS+SRBK) 45 238 5 13 6 307 5.20e5 12.83 225 3.98$\times$
Figure 12 109K, 80K, 200K GIPC 106 606 20 54 6 792 2.34e5 24.31 60 1.00$\times$
109K, 80K, 200K Ours (SRBK) 75 489 20 55 6 645 2.37e5 24.13 61 1.23$\times$
109K, 80K, 200K Ours (CEMAS+SRBK) 75 184 20 55 6 340 2.61e5 24.69 66 2.33$\times$
109K, 80K, 200K CIPC (CEMAS+SRBK) 104 202 21 91 6 424 2.89e5 26.25 69 -
100K, 0, 200K Ours (ABD+CEMAS+SRBK) 67 145 19 51 6 288 2.06e5 23.50 55 2.75$\times$
Figure 14 203K, 1122K, 50K GIPC 90 1481 4 14 7 1596 6.23e5 9.78 354 1.00$\times$
114K, 561K, 50K ZeMa 42 54810 2.7 9 7 54871 - 9.67 - 0.03$\times$
203K, 1122K, 50K Ours (SRBK) 63 569 3 14 7 656 6.23e5 9.78 354 2.43$\times$
203K, 1122K, 50K Ours (CEMAS+SRBK) 63 360 3 14 7 447 5.30e5 9.78 301 3.57$\times$
114K, 561K, 50K Ours (ABD+CEMAS+SRBK) 37 104 2.7 8.5 7 159 1.73e5 9.61 100 10.03$\times$
Figure 20 123K, 330K, 211K GIPC 137 3910 35 83 8 4173 4.31e6 39.35 548 1.00$\times$
123K, 330K, 211K Ours (SRBK) 91 2236 35 83 8 2453 4.38e6 39.65 552 1.70$\times$
123K, 330K, 211K Ours (CEMAS+SRBK) 90 730 35 82 8 945 1.81e6 39.25 231 4.42$\times$
Figure 21 79K, 298K, 110K GIPC 24 1080 5 13 8 1130 1.42e6 8.10 877 1.00×1.$0$
79K, 298K, 110K Ours (SRBK) 16 553 5 13 8 595 1.42e6 8.10 877 1.90$\times$
79K, 298K, 110K Ours (CEMAS+SRBK) 16 216 5 12 8 257 7.53e5 8.10 465 4.40$\times$
78K, 290K, 110K Ours (ABD+CEMAS+SRBK) 14 148 4 8 7 181 4.50e5 7.05 319 6.24$\times$
Figure 22 70K, 241K, 106K GIPC 692 10686 114 637 14 12143 7.30e6 15.80 462 1.00$\times$
70K, 241K, 106K Ours (SRBK) 460 4511 113 594 14 5692 7.58e6 15.80 480 2.13$\times$
70K, 241K, 106K Ours (CEMAS+SRBK) 452 1547 112 597 14 2722 3.37e6 15.70 214 4.46$\times$
53K, 133K, 106K Ours (ABD+CEMAS+SRBK) 407 696 103 535 12 1753 1.48e6 14.60 101 6.93$\times$
Figure 23 105K, 10K, 202K GIPC 210 2473 54 211 9 2957 3.10e6 37.72 329 1.00$\times$
105K, 10K, 202K Ours (SRBK) 142 762 51 205 9 1069 3.13e6 37.08 338 2.76$\times$
105K, 10K, 202K Ours (CEMAS+SRBK) 145 451 52 207 9 864 8.21e5 38.00 86 3.42$\times$
105K, 10K, 202K CIPC (CEMAS+SRBK) 219 812 59 329 10 1429 1.75e6 38.76 180 -
105K, 0, 202K Ours (ABD+CEMAS+SRBK) 161 468 47 178 9 863 7.78e5 34.04 91 3.42$\times$
@@
Here, we compare our method with GIPC (Huang et al., 2024) and ZeMa (Du et al., 2024) on the stiff and soft bunnies example (Figure 14). In this simulation, the Young’s modulus is set to $10^{6}$ Pa for the bottom bunny and $10^{8}$ Pa for the top bunny. Our framework can treat the top bunny as a stiff affine body and the bottom bunny with FEM, while GIPC needs to simulate both using FEM. Results in Table 3 indicate that our coupling framework achieves a 10$\times$ overall speedup compared to GIPC. To analyze the sources of this acceleration, we applied our contributions incrementally. ‘Ours (SRBK)’ uses only our SRBK method, matching other components to GIPC, which already yields a 2.43×2.43$\times$2.43 × speedup with more efficient Hessian assembly and SpMV. ‘Ours (CEMAS+SRBK)’ further adds connectivity-enhanced (CE) MAS preconditioner to ‘Ours (SRBK)’, retaining other components, and shows an additional 1.47×1.47$\times$1.47 × overall speedup due to faster PCG convergence enabled by CEMAS, resulting in a 3.57×3.57$\times$3.57 × improvement compared to GIPC. Note that in both ‘Ours (SRBK)’ and ‘Ours (CEMAS+SRBK)’ cases, we simulate both bunnies using FEM. ZeMa performs the worst in this case, as direct solvers are not efficient for simulations with a large number of DOFs. Thus, for general IPC simulations, GIPC remains the best alternative. Therefore, in subsequent experiments, we compare our methods only against GIPC.
@@
We further compare our method with GIPC on the examples shown in Figure 11, Figure 12, Figure 20, Figure 21, Figure 23, and Figure 22, with statistics reported in Table 3. Generally, ‘Ours (SRBK)’ achieves a 1.23×1.23$\times$1.23 ×2.76×2.76$\times$2.76 × speedup over GIPC. Adding our CEMAS preconditioner further improves performance, providing an additional 1.24×1.24$\times$1.24 ×2.6×2.6$\times$2.6 × speedup as shown by ‘Ours (CEMAS+SRBK)’, leading to an overall improvement of 2.33×2.$3$4.46×4.46$\times$4.46 × compared to GIPC. Notably, this is achieved without approximating stiff materials as affine bodies, which can further boost performance up to 2.75×2.75$\times$2.75 ×10×10$\times$10 × faster than GIPC. In Figure 12, comparing ‘Ours (SRBK)’ and ‘Ours (CEMAS+SRBK)’ reveals that while CEMAS does not improve PCG convergence due to similar ordering with Morton code sorting in GIPC, it significantly reduces PCG solver time due to lower construction and preconditioning costs. Here, PCG costs represent a smaller portion of total timing than in other examples due to fewer PCG iterations required, resulting in a modest 1.23×1.23$\times$1.23 × speedup with SRBK alone. We will discuss Figure 22 and Figure 23 in detail in the next subsection on stress tests.